OpenStudy (anonymous):

Find the exact value. sin-1(sqrt3/2)

OpenStudy (anonymous):

60 degrees... you can use the calculator to calculate it :)

OpenStudy (anonymous):

How would I type that into my calculator? Because when I try to, it says error.

OpenStudy (anonymous):

And 60 degrees is equal to pi/3, correct?

OpenStudy (anonymous):

forget degrees you are a grown up

OpenStudy (anonymous):

you are looking for the number between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose sine is \(\frac{\sqrt{3}}{2}\) that number is \(\frac{\pi}{3}\)

OpenStudy (anonymous):

Okay, but how do I figure out what number that is? I have no clue how to do this problem.

OpenStudy (anonymous):

you do it by knowing specific values for sine and cosine, and working backwards in other words, somewhere in your mind you know that \[\sin(\frac{\pi}{3})=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

or, you look at the second coordinate on the unit circle in the attached cheat sheet http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

when you see \(\frac{\sqrt{3}}{2}\) as the second coordinate, (which is sine) then look at the "angle" that gave it, you will see \(\frac{\pi}{3}\)

OpenStudy (anonymous):

Ok, printing that! Thanks.